A man raises a bucket of water of mass 10 kg from a well. When the bucket has risen 18 m, its velocity is 16 m/s. Work done by the man is (take g = 10 m/s2)


        (A)    3080 J


        (B)    1280 J


        (C)    1380 J


        (D)    1800 J



Grade:

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear ishika


work done by man = increse in total energy of bucket

                           =mgh + 1/2 m v2

                          = 10*10*18   + 1/2  * 10 * 162

                          = 3080 j


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Aditi Chauhan
14 years ago

 Dear ishika,

 

since

work done = potential energy+ kinetic energy.

                       = mgh + 1/2(mv2)

                        = (10×10×18)+[1⁄2×(162)].

                        = 3080 J

 

Thanks ans regards

Akhilesh Shukla